OpenStudy (anonymous):

How would I calculate the derivative of f(x)= (x^3+5)/x

OpenStudy (anonymous):

Please help me step by step (and explain why where you can) \[f(x)= \frac{ x ^{3}+5 }{ x }\] I have the answer, but I can't figure out how to get there: \[\frac{ 2x-5 }{ x ^{2} }\] Note: I'm supposed to be using the power rule.

OpenStudy (amistre64):

if you dont want to implement the usually convoluted quotient rule on a quotient; you might want to up the denominator and run teh product rule

OpenStudy (amistre64):

[(x^3+5) (x^-1) ]' = (x^3+5)' (x)+(x^3+5) (x^-1)' (3x^2) (x) - (x^3+5) (x^-2)

OpenStudy (anonymous):

I have to use the power rule on this one (as per the course requirements.) I'm just not sure how to go about it.

OpenStudy (amistre64):

you have to use more than a power rule .... you have a rational function, a quotient. you cant just use one single "rule"

OpenStudy (amistre64):

the "power rule" is required within it yes; but you cant just run that rule

OpenStudy (anonymous):

I'm allowed to pull the x up to the top temporarily, but I have to use the power rule and then I can bring the x back down.

OpenStudy (amistre64):

\[f(x)= \frac{ x ^{3}+5 }{ x }\] \[f'(x)= \frac{ x(x ^{3}+5)'-x'(x ^{3}+5) }{ x^2 }\]

OpenStudy (amistre64):

\[f(x)= \frac{ x ^{3}+5 }{ x }\] \[f(x)= (x ^{3}+5)~(x^{-1})\] \[f'(x)= (x ^{3}+5)'~(x^{-1})+(x ^{3}+5)~(x^{-1})'\]

OpenStudy (amistre64):

maybe expand it after pulling it up? \[f(x)= (x ^{3}+5)~(x^{-1})\] \[f(x)= x ^{2}+5x^{-1}\] \[f'(x)= 2x -5x^{-2}\]

OpenStudy (amistre64):

if so then its just playing with fractions again \[2x -5x^{-2}\] \[2x -\frac{5}{x^{2}}\] \[\frac{2xx^2}{x^2} -\frac{5}{x^{2}}\] \[\frac{2x^3-5}{x^2}\]